Answer the following questions regarding a 10x Serial Dilution experiment using yeast cell cultures from tube A.

Tube A contains 1 ml of yeast culture. 3.6 ml of sterile water is placed into 4 tubes (B, C, D, and E). 0.4 ml of yeast culture is removed from tube A and put into tube B.. 0.4 ml is transferred from tube B to tube C, and from tube C to tube E.

1.) What is the final volume in tubes C & E after the dilution?

C ____ml.   

E _____ ml

When 0.1 ml of Dilution D was plated on each of the four separate YED plates; 125, 140, 137, and 830 colonies were observed for each plate after 48 hours of incubation. When 0.1 ml of of dilution E was plated on each of the four sepearte YED plates; 12, 10, 11 and 8 colonies were observed on each plate after 48 hours of incubation.

2.) What is the cell denisity (titer) in tube D? (show your calculation and unit)

3.) What is the cell denisity (titer) in tube A? (show your calculation and unit)

4.) What is the cell denisity (titer) in tube A if the volume of cell culture plated on each plate was 50 microliters? (show your calculation and unit)
Answer #1

1. As C & E tubes contain 3.6 ml of sterile water. To C Tube, o.4 ml added from B. Again 0.4 from tube C taken out and added to E. So then C tube contain 3.6ml volume and E tube contain 4.0 ml volume after dilution

2. The titer of tube D = 1/1000: On an average D tube contain 3080 Bacteria/ml: (Average taken: 125+140+137+830/4 = 308/0.1ml or 3080/1ml)

3. A tube contain 3080 x 1000 = 3080000 Bacteria/ml (Because A tube is 1000 times concentrated than Tube-D). 1 ml contain 3080000 Bacteria. If 50 ul or 0.05 ml plated, then the cell density is 3080000/20

= 154000 Bacteria (Because 50 ul is 20 times less volume than original 1 ml volume of Tube - A

A -------1/10----------B----------1/100-------C-----------1/1000-----------D-----------1/10000-----------E